# arXiv:2603.01059v1[cs.CL] 01 Mar 2026

### 01 Mar 2026 v1

@@ 2. Related Work
Unlike single-user chatbots, limited research on group chatbots restricts their application in tasks like brainstorming sessions and debates. To bridge this gap, the MUCA framework (Mao et al., 2024) was the first to formalize the “3W” design dimensions—“What” to say, “When” to respond, and “Who” to answer—introducing conversational strategies for goal-oriented group discussions.  Building upon this foundation, the HUMA framework (Jacniacki and Serrat, 2025) further enables multi-user chatbots to achieve more natural and human-like behavior by leveraging human-like conversational strategies and response timing. Beyond conversational strategy design, other work has explored group chatbots from application and system-level perspectives. For example, MAP (Lee et al., 2025) proposes a multi-agent framework for multi-user personalization, through a reflection–analysis–feedback workflow to model user preferences, resolve conflicts, and iteratively refine personalized outcomes in collaborative settings. Similarly, Social-RAG (Wang et al., 2025) introduces a group agent for paper recommendations that retrieves social signals from group interactions to socially ground generation.  And SeeSawBot (Wang et al., 2026b) designs a multi-user chatbot to explore how LLM can mediate, facilitate, or disrupt group dynamics in collaborative environments. Besides, (Liu et al., 2025) has explored proactive agents with “inner thoughts” mechanisms and (Karahodža et al., 2025) proposes a conceptual framework that models group structures and interaction patterns.
@@ Appendix A Future Work
  • Group-level personalization. Future systems could move beyond individual personalization and model collective memory and shared concepts within a group. For example, a chatbot could remember entities (e.g., a user’s pet or personal images shared previously) and consistently recognize them in future multimodal interactions. This direction is closely related to recent advances in personalized agents (Woźniak et al., 2024; Hao et al., 2025; Alaluf et al., 2024; Nguyen et al., 2024), but extends personalization to the group level.
  • Synthetic data generation for group chats. Collecting real-world multi-user conversation data is costly and time-consuming. An important direction is to leverage LLMs and multimodal generative models (e.g., image and video generation) to synthesize high-fidelity group chat data that approximates real-world distributions, enabling both standalone and hybrid training paradigms. A promising line of work (Lee et al., 2021; Aboutalebi et al., 2024; Lee et al., 2024) explores transforming text-only dialogues into multimodal ones by inserting or replacing utterances with semantically aligned visual content, providing a practical pathway for scalable multimodal data construction. Another compelling approach involves role-playing with personalized agents. Recent studies (Xie et al., 2025; Wang et al., 2026a; Binz et al., 2025; Lei et al., 2026) have demonstrated that role-playing models, when conditioned on high-quality historical data of individuals, can simulate and predict human behavior in specific scenarios.  So, it becomes possible to generate more authentic and diverse group chat interactions that reflect real-world conversational dynamics.
@@
Appendix B Qualitative Analysis
@@
Appendix C Other materials
Additional materials will be provided in the next version of this paper, including detailed prompts, comprehensive descriptions of the MUIR dataset, and further system design details.

### 17 Mar 2026 v2

# GroupGPT: A Token-efficient and Privacy-preserving Agentic Framework for Multi-User Chat Assistant

@@ 2. Related Work
Unlike single-user chatbots, limited research on group chatbots restricts their application in tasks like brainstorming sessions and debates. To bridge this gap, the MUCA framework (Mao et al., 2024) was the first to formalize the “3W” design dimensions—“What” to say, “When” to respond, and “Who” to answer—introducing conversational strategies for goal-oriented group discussions.  Building upon this foundation, the HUMA framework (Jacniacki and Serrat, 2025) further enables multi-user chatbots to achieve more natural and human-like behavior by leveraging human-like conversational strategies and response timing. Beyond conversational strategy design, other work has explored group chatbots from application and system-level perspectives. For example, MAP (Lee et al., 2025) proposes a multi-agent framework for multi-user personalization, through a reflection–analysis–feedback workflow to model user preferences, resolve conflicts, and iteratively refine personalized outcomes in collaborative settings. Similarly, Social-RAG (Wang et al., 2025) introduces a group agent for paper recommendations that retrieves social signals from group interactions to socially ground generation. Besides, (Liu et al., 2025) has explored proactive agents with “inner thoughts” mechanisms and (Karahodža et al., 2025) proposes a conceptual framework that models group structures and interaction patterns.
@@ Appendix A Future Work
  • Group-level personalization. Future systems could move beyond individual personalization and model collective memory and shared concepts within a group. For example, a chatbot could remember entities (e.g., a user’s pet or personal images shared previously) and consistently recognize them in future multimodal interactions. This direction is closely related to recent advances in personalized agents (Woźniak et al., 2024; Hao et al., 2025; Alaluf et al., 2024; Nguyen et al., 2024; Jiang et al., 2025a, b), but extends personalization to the group level.
  • Synthetic data generation for group chats. Collecting real-world multi-user conversation data is costly and time-consuming. An important direction is to leverage LLMs and multimodal generative models (e.g., image and video generation) to synthesize high-fidelity group chat data that approximates real-world distributions, enabling both standalone and hybrid training paradigms. A promising line of work (Lee et al., 2021; Aboutalebi et al., 2024; Lee et al., 2024) explores transforming text-only dialogues into multimodal ones by inserting or replacing utterances with semantically aligned visual content, providing a practical pathway for scalable multimodal data construction. Another compelling approach involves role-playing with personalized agents. Recent studies (Xie et al., 2025; Wang et al., 2026; Binz et al., 2025; Lei et al., 2026) have demonstrated that role-playing models, when conditioned on high-quality historical data of individuals, can simulate and predict human behavior in specific scenarios. So, it becomes possible to generate more authentic and diverse group chat interactions that reflect real-world conversational dynamics.
@@
Appendix B Details of Dataset
In this section, we provide additional details of the MUIR dataset, including topic coverage, label distribution, intervention distance statistics, and the design considerations behind the data construction pipeline.
Figure 4. Word cloud visualization of frequently occurring discussion topics in the MUIR dataset.
B.1. Topic Diversity
To understand the diversity of conversations in MUIR, we analyze the topics appearing in the collected group chat segments. Figure 4 presents a word cloud visualization generated from the conversation messages in the dataset. Frequently occurring words correspond to common discussion themes such as daily life, technical questions, learning resources, entertainment, and coordination among group members. This visualization provides an intuitive overview of the semantic space covered by the dataset and indicates that MUIR captures a wide variety of natural group discussion topics.
B.2. Label Distribution
MUIR focuses on the intervention decision problem, where the assistant determines whether it should intervene in a group conversation. Each conversation segment is annotated with an intervention label indicating whether the assistant is expected to respond at that moment.
Figure 5 presents the distribution of intervention labels from two complementary perspectives. Specifically, Figure 5(b) reports the distribution of final intervention labels, which are computed solely from the training targets. By adjusting the parameter $X$ in Algorithm 4, we explicitly control the proportion of stay silent instances, such that it ultimately becomes the second most frequent label. This design encourages the model to learn a balanced intervention policy. In contrast, Figure 5(a) illustrates the distribution of all intervention labels, aggregated over both the conversation context and the final labels. This provides a more comprehensive view of intervention patterns in realistic group conversations. As shown in the Figure 5(a), emotional support and style balancing emerge as the most prevalent intervention types. This observation aligns well with real-world group chat dynamics, where providing emotional reassurance and resolving interpersonal tensions are among the most common and necessary forms of intervention.
Overall, the dataset captures a diverse yet realistic distribution of intervention behaviors, highlighting the importance of selective and context-aware participation by conversational agents.
(a) Distribution of all intervention labels in the MUIR dataset.
(b) Distribution of final intervention labels in the MUIR dataset.
Figure 5. Label distribution statistics of the MUIR dataset.
B.3. Intervention Distance Distribution
In real-world group conversations, assistant interventions do not occur uniformly but are typically separated by varying numbers of messages. To characterize this property, we measure the intervention distance, defined as the number of messages between two consecutive intervention labels.
Figure 6 illustrates the distribution of intervention distances in MUIR. The results show that most interventions occur after several messages rather than immediately following another intervention, reflecting the natural conversational dynamics in multi-user chat environments. Due to our annotation protocol, which intentionally adopts a relatively proactive intervention strategy, intervention instances are comparatively frequent. As a result, the average distance between two consecutive interventions is approximately 6 messages, with a median distance of 5 messages.
Figure 6. Distribution of message distances between consecutive intervention labels in the MUIR dataset.
B.4. Dataset Construction Pipeline
Constructing the MUIR dataset is a non-trivial process that involves several non-obvious design trade-offs. The primary objective of MUIR is to train a lightweight intervention judge model that can autonomously determine when to intervene and why, given a segment of group conversation. Formally, the model takes a sequence of recent messages as input and predicts an intervention decision for the latest message.
RQ1: How To Prevent Label Leakage in Sample Construction.
To construct MUIR, a straightforward strategy is to treat each intervention in the long group chat log as label and extract a fixed-size preceding context window as input. However, this naive approach introduces label leakage: intervention labels from one sample may appear in the context of another sample, allowing the model to trivially infer the target label from contextual cues rather than learning genuine decision patterns. Removing all historical interventions from the context is not viable, as it introduces additional issues discussed in RQ2.
Solution. To mitigate this issue, we adopt a sliding window strategy. Instead of using all intervention points as independent labels, only a subset of interventions is selected as prediction targets, while the remaining interventions are retained in the context as historical interventions. This design ensures that the model can learn from the context to fully exploit all annotated intervention signals in the training data, while avoiding label leakage across samples.
RQ2: What Should Be Included in Context?
Another critical question is what information should be exposed to the Intervention Judge. In particular, whether the context should include: (1) historical intervention reasons, and (2) responses generated by the final respondent. Including historical intervention reasons can help prevent redundant interventions triggered by the same cause within a short time span. Additionally, users may initiate follow-up interactions based on previous responses, which requires the model to recognize such scenarios and thus necessitates access to historical responses in the context.
We analyze several alternative designs:
Case 1: No historical interventions or responses. This setting leads to repeated or redundant interventions during inference. In many cases, the correct label should be stay silent, but the model may incorrectly trigger intervention because it cannot recognize that the same reason has already been addressed. Moreover, it fails to handle multi-turn user interactions with the GroupGPT.
Case 2: Include responses but exclude historical interventions. Observing responses from the final respondent implies that the model must wait for the LLM’s output during inference, introducing additional latency. Furthermore, responses are strongly correlated with intervention reasons, which leads to implicit label leakage.
Case 3: Include both responses and historical interventions. This setting inherits the drawbacks of increased inference latency and potential data leakage. In addition, it requires realistic multi-turn interaction data between users and the group chat agent, which is not sufficiently available in practice.
Final Design.
Our final pipeline integrates the considerations from both RQ1 and RQ2:
  • We employ a sliding window mechanism, where only a subset of interventions is used as prediction labels, while the remaining interventions are preserved as historical context. This allows the model to learn intervention patterns without label leakage.
  • We exclude historical responses from the context to reduce inference latency.
  • To handle user-initiated multi-turn interactions, we introduce an explicit protocol: when users directly address the agent (e.g., via @groupgpt), the intervention judge is bypassed, and the request is forwarded to the final respondent for response generation through careful prompt design.
This design achieves a balance between learning effectiveness, inference efficiency, and robustness in real-world conversational settings.
Appendix C Prompts
The detailed prompts used in our experiments are provided in the code repository.
Appendix D Qualitative Analysis